OpenStudy (jwest6894):

Help, i kind of know it but forgot how to do it exactly. Simplify each expression. Express the answer so that all exponents are positive. Help, i kind of know it but forgot how to do it exactly. (3x^2/5y^4)^2

OpenStudy (calculusxy):

\[(\frac{3x^2}{5y^2})^2 = \frac{3^2x^4}{5^2y^4} = \frac{9x^4}{25y^4} \]

OpenStudy (jwest6894):

how did u do that first step

OpenStudy (calculusxy):

It's called the power rule. It's when you distribute the exponent that is outside of the parenthesis to the terms that are in the parenthesis. Therefore ,I just distributed the 2 to 3 and got \(3^2\) and then to \(x^2\) to get \(x^{2 \times 2} = x^4\). The same thing was done to the denominator too.

OpenStudy (calculusxy):

Do you understand it?

OpenStudy (jwest6894):

OK i got it... sorry for late reply.. had to do something, but thanks